# Projective reflection groups of finite covolume

### v1

@@ 3.3. Cartan matrices
If $A$ is a Cartan matrix on a finite set $S$, then for any subset $T$ of $S$, we let $A_{T}\coloneqq(A_{st})_{s,t\in T}$. It is also a Cartan matrix. Moreover, observe that the Coxeter group $W_{A}$ allows to define the relation $\perp$ on $S$, so we can define as before the irreducible components of $S$. In an ordering of $S$ compatible with the decomposition of $S$ into its irreducible components $S_{1}\sqcup\dots\sqcup S_{k}$, we see that $A$ splits into a block diagonal matrix whose blocks are the Cartan matrices $A_{S_{i}}$. We call these Cartan submatrices of $A$ the irreducible components of $A$.
A Cartan matrix $A$ is irreducible if it has a single irreducible component, or equivalently, if the associated Coxeter group $W_{A}$ is irreducible.
If $A$ is an irreducible Cartan matrix, then 2IA2I-A is irreducible with non-negative entries, so by the Perron–Frobenius theorem, its complex eigenvalue of largest modulus is real, positive and simple. In other words, $A$ admits a real simple eigenvalue $\lambda_{A}$ such that $|2-\mu|$ is maximal at $\lambda_{A}$, where $\mu$ is any complex eigenvalue of $A$.
@@
If $A$ is reducible, we say it is of positive (resp. zero, resp. negative) type if all the irreducible component of $A$ are of positive (resp. zero, resp. negative) type.
Theorem 3.11 ([Vin71, Thm. 3]).
@@
Let $A$ be a (non-necessarily irreducible) Cartan matrix on $S$ of positive (resp. zero, resp. negative) type. Then there is a vector $X>0$ in $\mathbb{R}^{S}$ such that $AX>0$ (resp. $AX=0$, resp. $AX<0$). That is because we can split $S$ into its irreducible components $S_{1},\dots,S_{k}$ and choose for each $i$ a vector $X_{i}>0$ in $\mathbb{R}^{S_{i}}$ such that $A_{S_{i}}X_{i}$ has the right sign, by Perron–Frobenius theorem. Since $A$ is the block diagonal matrix with blocks $(A_{S_{i}})_{i=1}^{k}$, we can then choose $X=(X_{s})_{s\in S}$ where $X_{s}=(X_{i})_{s}$ if $s\in S_{i}$.
@@ 3.4. Reduced and dual-reduced Vinberg representations
$V_{\alpha}$ $\coloneqq\bigcap_{s\in S}\ker\alpha_{s}$ $V_{v}$ $\coloneqq\operatorname{Span}\{v_{s}\mid s\in S\}$ $V_{v}^{\alpha}$ $\coloneqq V_{v}/(V_{\alpha}\cap V_{v}).$
@@
The following was proven in [Vin71] (see Equation (32) and the corolloray to Proposition 19). The addition when $W$ is large is from [Mar17, Thm. 2.18].
Proposition 3.13.
@@
Proposition 3.14 ([DGKLM25, Cor. 3.25]).
Let $\rho:W\to\operatorname{SL^{\pm}}(V)$ be a Vinberg representation with a Cartan matrix of negative type. Then $\Lambda_{\rho(W)}$ is non-empty and lies in $\operatorname{\mathbb{P}}(V_{v})$.
@@ 4. Coxeter polytopes
Since the data of a Coxeter polytope $P$ is enough to define each reflections $(\sigma_{s})_{s\in S}$, there is an associated Cartan matrix $A_{P}$, Coxeter group $W_{P}$, reflection group $\Gamma_{P}$, and Vinberg representation $\rho_{P}:W_{P}\to\operatorname{SL^{\pm}}(V)$. By Theorem 3.8, $\Gamma_{P}$ preserves an open convex cone $C_{P}$ tiled by $C_{P}\cap\Delta$, so the Vinberg domain
@@ 4.1. Types of Coxeter polytopes
Similarly, we say that $P$ is irreducible if $A_{P}$ (or equivalently, $W_{P}$) is irreducible.
If $A_{P}$ is of positive type, then we have $\operatorname{rk}A_{P}=d+1$ by Theorem 3.11. One can produce examples of non-parabolic Coxeter polytopes of zero type and non-loxodromic Coxeter polytopes of negative type by Remark 3.9 using examples of such Cartan matrices.
@@
Proposition 4.5 ([Mar17, Prop. 2.9]).
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a Coxeter polytope. If $P$ is of negative type, either $W_{P}$ is affine of type $\tilde{A}_{d}$, or it is large. If $W_{P}$ is large, then $P$ is of negative type.
@@ 4.2. Indecomposable Coxeter polytopes
Let $C_{1}\subset V_{1}$ and $C_{2}\subset V_{2}$ be two sharp open convex cones in the finite dimensional real vector spaces $V_{1}$ and $V_{2}$. The sum $C_{1}+C_{2}$ is a sharp open convex cone of $V_{1}\oplus V_{2}$. Let now $\Omega_{1}\coloneqq\operatorname{\mathbb{P}}(C_{1})$ and $\Omega_{2}\coloneqq\operatorname{\mathbb{P}}(C_{2})$. A join of $\Omega_{1}$ and $\Omega_{2}$ is a properly convex domain of $\operatorname{\mathbb{P}}(V_{1}\oplus V_{2})$ of the form $\operatorname{\mathbb{P}}(C_{1}+C_{2})$. Notice that since there are two different sharp open convex cones lifting $\Omega_{1}$ and $\Omega_{2}$ (the cones and their opposite), there are $2$ different joins. More generally, there are $2^{r-1}$ different joins of $r$ properly convex domains. Observe that all the joins are image of one another by an element of $\operatorname{SL^{\pm}}(V)$, so it doesn’t usually makes much difference to work with one of them rather than the other. Moreover, since the subgroup of $\operatorname{SL^{\pm}}(V_{1})$ preserving $C_{1}$ also preserves $C_{1}$, and the same holds for $C_{2}$, it is clear that all the joins of $\Omega_{1}$ and $\Omega_{2}$ are invariant under exactly the same elements of $\operatorname{SL^{\pm}}(V)$.
@@
Note that an irreducible Coxeter polytope is necessarily indecomposable.
@@ 4.3. Links of Coxeter polytopes
Let $P$ be a Coxeter polytope and $f$ be a face of $P$. Recall from Theorem 3.8 that the stabilizer of $f$ in $\Gamma_{P}$ is $\Gamma_{f}\coloneqq\rho_{P}(W_{S_{f}})$ where $S_{f}=\{s\in S\mid f\subset\operatorname{\mathbb{P}}(\ker\alpha_{s})\}$. The restriction of $\rho_{P}$ to $W_{S_{f}}$ is a Vinberg representation with Cartan matrix $A_{S_{f}}$ which is not reduced as $\bigcap_{s\in S_{f}}\ker\alpha_{s}=\operatorname{Span}f$ is non-trivial. However, this representations induces a reduced Vinberg representation $W_{S_{f}}\to\operatorname{SL^{\pm}}(V)$. The associated Coxeter polytope is the link $P_{f}\subset\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ of $P$ at $f$, given by the projection in $\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ of the convex cone $\bigcap_{s\in S_{f}}\{\alpha_{s}\leq 0\}$.
@@
Let $P$ be a Coxeter polytope. By Theorem 3.8, one can see that $\Gamma_{P}$ acts cocompactly on $\Omega_{P}$ if and only if $P\subset\Omega_{P}$, if and only if all its vertices lie in $\Omega_{P}$, if and only if all its vertices are elliptic. This motivates the following definition.
@@
Elliptic and parabolic Coxeter polytopes are always perfect by [Vin71, Prop. 11]. Vinberg also weakened slightly the notion of a perfect Coxeter polytope in order to accommodate the reflections groups in $\mathbb{H}^{d}$ that act with finite covolume.
@@ 4.4. Maximality of the Vinberg domain
Proposition 4.10 ([DGKLM25, Prop. 4.1]).
Let $P\subset\operatorname{\mathbb{P}}(V)$ be an irreducible Coxeter polytope of negative type such that $|S|\geq 3$. Then $\Omega_{P}$ is the largest $\Gamma$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$. In other words, $\Omega_{P}$ contains any $\Gamma$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$.
Danciger–Guéritaud–Kassel–Lee–Marquis also give a description of the smallest invariant properly convex domain defined in Proposition 2.6.
Proposition 4.11 ([DGKLM25, Thm. 5.2]).
Suppose $P\subset\operatorname{\mathbb{P}}(V)$ is an irreducible Coxeter polytope of negative type such that $\operatorname{Span}(v_{s})_{s\in S}=V$. Then
@@
is the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$ contained in $\Omega_{P}$. In other words, $\Omega_{\rm min}$ is contained in all $\Gamma$-invariant properly convex domains of $\operatorname{\mathbb{P}}(V)$ contained in $\Omega_{P}$.
@@
In particular, if $\Gamma_{P}$ is strongly irreducible, then $\Omega_{\text{min}}$ also equals $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ and is the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$ by Proposition 2.6. This holds in particular when $\Gamma_{P}$ is a large Coxeter group, by Proposition 3.13.
@@ 5.1. No proper faces of negative type
Let $V\coloneqq E_{1}\oplus E_{2}$ be a non-trivial decomposition, and $\Omega_{i}\subset\operatorname{\mathbb{P}}(E_{i})$ be properly convex domains for $(i=1,2)$. Let $\Omega\coloneqq\Omega_{1}\mathbin{\otimes}\Omega_{2}$ be a join of $\Omega_{1}$ and $\Omega_{2}$. If $P\subset\Omega$ is convex with nonempty interior, and $\Omega_{1}\cap\overline{P}$ has nonempty relative interior in $\Omega_{1}$, then $\mu_{\Omega}(P)=\infty$.
@@
Let $\pi:\operatorname{\mathbb{P}}(V)\smallsetminus\operatorname{\mathbb{P}}(E_{1})\to\operatorname{\mathbb{P}}(E_{2})$ be the projectivization of the projection onto $E_{2}$ parallel to $E_{1}$, and $Q_{1}\subset\operatorname{int}(\Omega_{1}\cap\overline{P})$ and $Q_{2}\subset\operatorname{int}\pi(P)$ be closed subsets of non-empty interior. Lastly, let $Q\coloneqq P\cap\operatorname{Conv}(Q_{1}\cup Q_{2})$ and $h\in\operatorname{GL}(V)$ defined by
@@
for all $(v_{1},v_{2})\in E_{1}\times E_{2}$. Observe that $h$ defines an automorphism of $\Omega$ that preserves all projective lines through a point of $\operatorname{\mathbb{P}}(E_{1})$ and a point of $\operatorname{\mathbb{P}}(E_{2})$, moving points on these lines closer to $\operatorname{\mathbb{P}}(E_{1})$. Since $Q$ is foliated by intervals of such projective lines with one endpoint on $Q_{1}\subset\operatorname{\mathbb{P}}(E_{1})$, we have $h(Q)\subset Q$. We may suppose up to shrinking $P$ that $\overline{P}\cap\operatorname{\mathbb{P}}(E_{2})=\varnothing$, in which case $Q\smallsetminus h(Q)$ contains a neighborhood of the union of the other endpoint of these intervals, so it has non-empty interior. It follows that $\mu_{\Omega}(P)=\mu_{\Omega}(hP)+\mu_{\Omega}(hP\smallsetminus P)>\mu_{\Omega}(P)$, so we have $\mu_{\Omega}(P)=+\infty$. ∎
Now Lemma 5.1 implies the following:
@@
Suppose $P$ has a face $f$ of negative type. The Vinberg domain $\Omega_{f}$ of the link $P_{f}$ is a properly convex domain of $\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ by Proposition 4.3. Let $E$ be a supplement of $\operatorname{Span}f$ in $V$ such that $\operatorname{\mathbb{P}}(E)$ does not meet the closure of $\Omega_{P}$. The decomposition $V=\operatorname{Span}f\oplus E$ allows to define a map $\phi:\operatorname{\mathbb{P}}(V/\operatorname{Span}f)\to\operatorname{\mathbb{P}}(E)$. Choose a properly convex domain $\Omega^{\prime}\subset\operatorname{\mathbb{P}}(\operatorname{Span}f)$ such that $\Omega^{\prime}$ contains $f$. Now if we let $\Omega\coloneqq\Omega^{\prime}\mathbin{\otimes}\phi(\Omega_{f})$, $\Omega$ contains $\Omega_{P}$, so Proposition 2.4 and Lemma 5.1 yield $\mu_{\Omega_{P}}(P)\geq\mu_{\Omega}(P)=+\infty$. ∎

### v2

@@ 3.3. Cartan matrices
If $A$ is a Cartan matrix on a finite set $S$, then for any subset $T$ of $S$, we let $A_{T}\coloneqq(A_{st})_{s,t\in T}$. It is also a Cartan matrix. Moreover, the Coxeter group $W_{A}$ allows to define the relation $\perp$ on $S$, so we can define as before the irreducible components of $S$. In an ordering of $S$ compatible with the decomposition of $S$ into its irreducible components $S_{1}\sqcup\dots\sqcup S_{k}$, the matrix $A$ splits into a block diagonal matrix whose blocks are the Cartan matrices $A_{S_{i}}$. We call these Cartan submatrices of $A$ the irreducible components of $A$.
A Cartan matrix $A$ is irreducible if it has a single irreducible component, or equivalently, if the associated Coxeter group $W_{A}$ is irreducible. If $A$ is an irreducible Cartan matrix, then 2IA2I-A is irreducible with nonnegative entries, so by the Perron–Frobenius theorem, its complex eigenvalue of largest modulus is real, positive and simple. In other words, $A$ admits a real simple eigenvalue $\lambda_{A}$ such that $|2-\mu|$ is maximal at $2-\lambda_{A}$, where $\mu$ is any complex eigenvalue of $A$.
@@
If $A$ is reducible (i.e. not irreducible), we say it is of positive (resp. zero, resp. negative) type if all the irreducible components of $A$ are of positive (resp. zero, resp. negative) type.
Fact 3.11 ([Vin71, Thm. 3]).
@@
Let $A$ be a (non necessarily irreducible) Cartan matrix on $S$ of positive (resp. zero, resp. negative) type. Then there is a vector $X>0$ in $\mathbb{R}^{S}$ such that $AX>0$ (resp. $AX=0$, resp. $AX<0$). That is because we can split $S$ into its irreducible components $S_{1},\dots,S_{k}$ and choose for each $i$ a vector $X_{i}>0$ in $\mathbb{R}^{S_{i}}$ such that $A_{S_{i}}X_{i}$ has the right sign, by Perron–Frobenius theorem. Since $A$ is the block diagonal matrix with blocks $(A_{S_{i}})_{i=1}^{k}$, we can then choose $X=(X_{s})_{s\in S}$ where $X_{s}=(X_{i})_{s}$ if $s\in S_{i}$.
@@ 3.4. Reduced and dual-reduced Vinberg representations
$V_{\alpha}$ $\coloneqq\bigcap_{s\in S}\ker\alpha_{s}$ $V_{v}$ $\coloneqq\operatorname{Span}\{v_{s}\mid s\in S\}$ $V_{v}^{\alpha}$ $\coloneqq V_{v}/(V_{\alpha}\cap V_{v})$
@@
The following was proven in [Vin71] (see Equation (32) and the corollary to Proposition 19). The addition when $W$ is large is from [Mar17, Thm. 2.18].
Fact 3.13.
@@
Fact 3.14 ([DGKLM25, Cor. 3.25]).
Let $\rho:W\to\operatorname{SL^{\pm}}(V)$ be a Vinberg representation with a Cartan matrix of negative type. Then $\rho(W)$ is proximal and $\Lambda_{\rho(W)}\subset\operatorname{\mathbb{P}}(V_{v})$.
@@ 4. Coxeter polytopes
Since the data of a Coxeter polytope $P$ is enough to define the reflections $(\sigma_{s})_{s\in S}$, there is an associated Cartan matrix $A_{P}$, Coxeter group $W_{P}$, reflection group $\Gamma_{P}$, and Vinberg representation $\rho_{P}:W_{P}\to\operatorname{SL^{\pm}}(V)$. By Fact 3.8, $\Gamma_{P}$ preserves an open convex cone $C_{P}$ tiled by $C_{P}\cap\Delta$, so the Vinberg domain
@@ 4.1. Types of Coxeter polytopes
Similarly, we say that $P$ is irreducible or reducible if $A_{P}$ (or equivalently, $W_{P}$) is so.
If $A_{P}$ is of positive type, then we have $\operatorname{rk}A_{P}=d+1$ by Fact 3.11. One can produce examples of non-parabolic Coxeter polytopes of zero type and non-loxodromic Coxeter polytopes of negative type by Remark 3.9 using examples of such Cartan matrices.
@@
Fact 4.5 ([Mar17, Prop. 2.9]).
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a Coxeter polytope. If $P$ is of negative type, either $W_{P}$ is affine of type $\tilde{A}_{d}$, or it is large.
If $W_{P}$ is large, then $P$ is of negative type.
@@ 4.2. Indecomposable Coxeter polytopes
Let $C_{1}\subset V_{1}$ and $C_{2}\subset V_{2}$ be two sharp open convex cones in the finite-dimensional real vector spaces $V_{1}$ and $V_{2}$. The sum $C_{1}+C_{2}$ is a sharp open convex cone of $V_{1}\oplus V_{2}$. Let now $\Omega_{1}\coloneqq\operatorname{\mathbb{P}}(C_{1})$ and $\Omega_{2}\coloneqq\operatorname{\mathbb{P}}(C_{2})$. A join of $\Omega_{1}$ and $\Omega_{2}$ is a properly convex domain of $\operatorname{\mathbb{P}}(V_{1}\oplus V_{2})$ of the form $\operatorname{\mathbb{P}}(C_{1}+C_{2})$. Notice that since there are two different sharp open convex cones lifting $\Omega_{1}$ and $\Omega_{2}$ (the cones and their opposite), there are $2$ different joins. More generally, there are $2^{r-1}$ different joins of $r$ properly convex domains. Observe that all the joins are image of one another by elements of $\operatorname{SL^{\pm}}(V)$, so it does not usually make much difference to work with one of them rather than the others. Moreover, since the subgroup of $\operatorname{SL^{\pm}}(V_{1})$ preserving $C_{1}$ also preserves $C_{1}$ and the same holds for $C_{2}$, it is clear that all the joins of $\Omega_{1}$ and $\Omega_{2}$ are invariant under exactly the same elements of $\operatorname{SL^{\pm}}(V)$.
@@
Note that an irreducible Coxeter polytope is necessarily indecomposable. It will be useful to describe the proximal limit set of a join of Coxeter polytopes.
Lemma 4.7.
Let $P$ be a Coxeter polytope that is a join of Coxeter polytopes $(P_{i})_{i=1}^{k}$ of negative type. Then $\Gamma_{P}$ is proximal and
$\Lambda_{P}=\bigcup_{i=1}^{k}\Lambda_{P_{i}}$
Proof.
Let $V=\bigoplus_{i=1}^{k}V_{i}$ be the decomposition of $V$ so that each $P_{i}$ is a Coxeter polytope in $\operatorname{\mathbb{P}}(V_{i})$. Observe that $\Gamma_{P}$ is isomorphic to $\Gamma_{P_{1}}\times\dots\times\Gamma_{P_{k}}$, with $\Gamma_{P_{i}}$ acting naturally on $\Omega_{P_{i}}$ and trivially on the other factors $\Omega_{P_{j}}$ for $j\neq i$. In particular, a proximal element in $\Gamma_{P}$ is of the form $(\gamma_{1},\dots,\gamma_{k})$ where $\gamma_{i}\in\Gamma_{P_{i}}$ is proximal for some $i$ with top eigenvalue larger than the modulus of the complex eigenvalues of the $\gamma_{j}$ for $j\neq i$. There exist proximal elements in $\Gamma_{P}$ as each $\Gamma_{P_{i}}$ is proximal by Fact 3.14, so $\Gamma_{P}$ is proximal. Moreover, by the description of the proximal elements of $\Gamma_{P}$ above, $\Lambda_{P}=\bigcup_{i=1}^{k}\Lambda_{P_{i}}$. ∎
@@ 4.3. Links of Coxeter polytopes
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a Coxeter polytope and $f$ be a face of $P$. Recall from Fact 3.8 that the stabilizer of $f$ in $\Gamma_{P}$ is $\Gamma_{f}\coloneqq\rho_{P}(W_{S_{f}})$ where $S_{f}=\{s\in S\mid f\subset\operatorname{\mathbb{P}}(\ker\alpha_{s})\}$. The restriction of $\rho_{P}$ to $W_{S_{f}}$ is a Vinberg representation with Cartan matrix $A_{S_{f}}$, which is not reduced as $\bigcap_{s\in S_{f}}\ker\alpha_{s}=\operatorname{Span}f$ is non-trivial. However, this representation induces a reduced Vinberg representation $W_{S_{f}}\to\operatorname{SL^{\pm}}(V/\operatorname{Span}f)$. The associated Coxeter polytope is the link $P_{f}\subset\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ of $P$ at $f$, given by the projection in $\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ of the convex cone $\bigcap_{s\in S_{f}}\{\alpha_{s}\leq 0\}$.
@@
Let $P$ be a Coxeter polytope. By Fact 3.8, one can see that $\Gamma_{P}$ acts cocompactly on $\Omega_{P}$ if and only if $P\subset\Omega_{P}$, if and only if all its vertices lie in $\Omega_{P}$, if and only if all its vertices are elliptic. This motivates the following definition.
@@
Elliptic and parabolic Coxeter polytopes are always perfect by [Vin71, Prop. 11]. Vinberg also slightly weakened the notion of a perfect Coxeter polytope in order to accommodate the reflections groups in $\mathbb{H}^{d}$ that act with finite covolume.
@@ 4.4. Maximality of the Vinberg domain
Fact 4.11 ([DGKLM25, Prop. 4.1]).
Let $P\subset\operatorname{\mathbb{P}}(V)$ be an irreducible Coxeter polytope of negative type such that $|S|\geq 3$. Then $\Omega_{P}$ is the largest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$. In other words, $\Omega_{P}$ contains any $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$.
Danciger–Guéritaud–Kassel–Lee–Marquis also gave a description of the smallest invariant properly convex domain defined in Fact 2.6.
Fact 4.12 ([DGKLM25, Thm. 5.2]).
Suppose $P\subset\operatorname{\mathbb{P}}(V)$ is an irreducible Coxeter polytope of negative type such that $V_{v}=V$. Then
@@
is the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$ contained in $\Omega_{P}$. In other words, $\Omega_{\rm min}$ is contained in all $\Gamma_{P}$-invariant properly convex domains of $\operatorname{\mathbb{P}}(V)$ contained in $\Omega_{P}$.
@@
In particular, if $\Gamma_{P}$ is strongly irreducible, then $\Omega_{\text{min}}$ also equals $\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ and is the smallest $\Gamma_{P}$-invariant properly convex domain of $\operatorname{\mathbb{P}}(V)$ by Fact 2.6. This holds in particular when $\Gamma_{P}$ is a large Coxeter group, by Fact 3.13.
@@ 5.1. No proper faces of negative type
Let $V\coloneqq E_{1}\oplus E_{2}$ be a non-trivial decomposition, and $\Omega_{i}\subset\operatorname{\mathbb{P}}(E_{i})$ be properly convex domains for $(i=1,2)$. Let $\Omega\coloneqq\Omega_{1}\mathbin{\otimes}\Omega_{2}$ be a join of $\Omega_{1}$ and $\Omega_{2}$. If $P\subset\Omega$ is convex with non-empty interior, and $\Omega_{1}\cap\overline{P}$ has non-empty interior in $\Omega_{1}$, then $\mu_{\Omega}(P)=\infty$.
@@
Let $\pi:\operatorname{\mathbb{P}}(V)\smallsetminus\operatorname{\mathbb{P}}(E_{1})\to\operatorname{\mathbb{P}}(E_{2})$ be the projectivization of the projection onto $E_{2}$ parallel to $E_{1}$, and $Q_{1}\subset\operatorname{int_{\text{rel}}}(\Omega_{1}\cap\overline{P})$ and $Q_{2}\subset\operatorname{int_{\text{rel}}}\pi(P)$ be closed subsets of non-empty interior. Lastly, let $Q\coloneqq P\cap\operatorname{Conv}(Q_{1}\cup Q_{2})$, and define $h\in\operatorname{GL}(V)$ by
@@
for all $(v_{1},v_{2})\in V=E_{1}\oplus E_{2}$. Observe that $\frac{1}{\lvert\det h\rvert}h$ defines an automorphism of $\Omega$ that preserves all projective lines through a point of $\operatorname{\mathbb{P}}(E_{1})$ and a point of $\operatorname{\mathbb{P}}(E_{2})$, moving points on these lines closer to $\operatorname{\mathbb{P}}(E_{1})$. Since $Q$ is foliated by intervals of such projective lines with one endpoint on $Q_{1}\subset\operatorname{\mathbb{P}}(E_{1})$, we have $h(Q)\subset Q$. By shrinking $P$ if necessary, we may suppose that $\overline{P}\cap\operatorname{\mathbb{P}}(E_{2})=\varnothing$. In this case, $Q\smallsetminus h(Q)$ contains a neighborhood of the union of the other endpoint of these intervals, so it has non-empty interior. It follows that $\mu_{\Omega}(Q)=\mu_{\Omega}(h(Q))+\mu_{\Omega}(Q\smallsetminus h(Q))>\mu_{\Omega}(Q)$, so we have $\mu_{\Omega}(P)\geq\mu_{\Omega}(Q)=+\infty$. ∎
By Lemma 5.1, we have
@@
Suppose $P$ has a proper face $f$ of negative type. The Vinberg domain $\Omega_{f}$ of the link $P_{f}$ is a properly convex domain of $\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ by Fact 4.3. Let $E$ be a supplement of $\operatorname{Span}f$ in $V$ such that $\operatorname{\mathbb{P}}(E)$ does not meet the closure of $\Omega_{P}$. We have an isomorphism $\phi:\operatorname{\mathbb{P}}(V/\operatorname{Span}f)\to\operatorname{\mathbb{P}}(E)$. Choose a properly convex domain $\Omega^{\prime}\subset\operatorname{\mathbb{P}}(\operatorname{Span}f)$ such that $\Omega^{\prime}$ contains $f$. Now if we let $\Omega\coloneqq\Omega^{\prime}\mathbin{\otimes}\phi(\Omega_{f})$, then $\Omega$ contains $\Omega_{P}$, and Fact 2.4 and Lemma 5.1 yield $\mu_{\Omega_{P}}(P)\geq\mu_{\Omega}(P)=+\infty$. ∎